OpenStudy (anonymous):

y-10=2(x-4) find x and y intercepts full work shown thanks :)

OpenStudy (anonymous):

put x=0, we can get the y-intercept which is -8+10=... put y=0, we can get the x-intercept which is -10/2 +4=-5+4=....

OpenStudy (anonymous):

y-10=2x-8 y=2x-8+10 y=2x+2 x intercept y=0 thus, 0=2x+2 x=-1 y intercept x=0 thus, y=2*0+2 y=2

mathslover (mathslover):

Remember whenever we do these type of questions : If we are going to find : Y intercept : put x = 0 and solve for y If we are going to find : X intercept : put y = 0 and solve for x best of luck